One of two supplementary angles is 4 degree more than one-third of the other angle. find the measure of each angle.
Answer by checkley77(12844)

SUPPLEMENTARY ANGLES (2) ADD UP YO 180 DEGREES.
THEREFORE:
X=Y/3+4 & X+Y=180
THUS WE HAVE THE EQUATION
(Y/3+4)+Y=180
Y/3+4+Y=180
Y/3+Y=180-4
(Y+3Y)/3=176
4Y=176*3
4Y=528
Y=528/4
Y=132 DEGREES IS THE LARGER ANGLE.
X=132=180
X=180-132
X=48 DEGREES FOR THE SMALLER ANGLE.
PROOF:
132/3+4+132=180
44+4+132=180
180=180